# Patent-CR: A Dataset for Patent Claim Revision

### v1

@@ 1 Introduction
Text revision aims to improve text quality, such as fixing grammar errors (Fang et al., 2023) and enhancing sentence coherence (Geva et al., 2019). Currently, datasets for this task are derived from scientific literature, Wikipedia entries, and news articles (Du et al., 2022). In this paper, we broaden the scope of text revision to encompass the domain of patents, characterized by large-scale, complex, and precise textual data. The patent domain presents unique opportunities and challenges for the field of artificial intelligence (AI) and natural language processing (NLP) (Jiang and Goetz, 2024).
@@
Patent claims are critical in a patent application document. As the legal centerpiece, they define the technical scope of the invention and ensure the patent can withstand legal scrutiny. We introduce the relevant background information of patents in Appendix A. Drafting and revising patent applications are both time-intensive and financially burdensome (LLP, 2023). Research showed that large language models (LLMs) have the potential to generate high-quality patent claims but current performance are not yet satisfactory (Jiang et al., 2024b). To facilitate the automation of patent writing, we propose a new task, namely patent claim revision, aiming at improving the quality of patent claims to pass the legal scrutiny of patent offices.
Figure 1 illustrates an example of patent claim revision. By checking the dataset and consulting patent professionals, we have identified five different types of modifications between the draft and final versions. (1) Content amendment: Essential information missing in the draft is included or unnecessary information is removed. (2) Term consistency: Technical terms are ensured to be consistent throughout the document. (3) Language precision: Grammatical errors are corrected, and word choice is refined for greater precision. (4) Concision: Some claims are merged with others for concision. (5) Renumbering: The consolidation of claims necessitates adjustments in their numbering.
@@ 2.2 Patent Writing
The adoption of LLMs in generating patent text primarily aims to enhance the efficiency and efficacy of drafting patent applications. Despite the potential capabilities of LLMs, current research in this area remains limited and largely unsatisfactory (Jiang and Goetz, 2024). An early study by Lee and Hsiang (2020) served as a preliminary exploration into generating patent claims with the fine-tuning of GPT-2 (Radford et al., 2019). The authors demonstrated that minimal training steps were adequate for the model to generate patent-like texts, but they did not evaluate the quality of the generated text. Subsequent research by Lee (2020) expanded on this aspect by training GPT-2 to convert one element of a patent application into another, for example, creating abstracts from titles and claims from abstracts. As the abstract is typically rather generic and imprecise, the latter may not be a well-conditioned task. Hence, Jiang et al. (2024b) proposed the description-based claim generation task and evaluated the performance of the current LLMs on this domain-specific task. We extend the task to claim revision to explore whether LLMs can further improve the quality of claims. Moreover, Christofidellis et al. (2022) introduced a prompt-based generative transformer (PGT) for patent-related tasks, which used GPT-2 as a foundational model and employed multi-task learning (MTL) (Maurer et al., 2016) to train on various tasks, including part-of-patent generation, text infilling, and evaluating patent coherence. Additionally, Aubakirova et al. (2023) presented the first large-scale patent figure-caption dataset, designed for patent figure caption generation.
@@ 3.1 Data Collection
Step 2: The European Patent Office provides the Open Patent Services (OPS) for public access to their data. [^3] https://www.epo.org/en/searching-for-patents/data/web-services/ops We retrieve the application and published versions of claims corresponding to specific patent publication numbers through the OPS API. A patent has different versions published by EPO, where A1 or A2 is the patent application and B1 is the granted patent. We primarily used the A1 version, and we used A2 if A1 is not available. We eliminated the patents that do not have either A1 or A2. We opted for the B1 version as the revised claims and discarded those without the B1 version.
Step 3: In the final compilation of our dataset, we formulate data into an easy-readable format and manually check in detail to ensure dataset’s quality.
@@ 4.1 Models
We use the original claims without any editing as a baseline similar to previous text revision works (Raheja et al., 2023; Jourdan et al., 2024), namely the Copy baseline. This baseline performs well when there are extensive overlaps between source inputs and target outputs. We can also evaluate other models’ performance by comparing the results with this simple baseline. We select the state-of-the-art CoEdIT-XL (Raheja et al., 2023) as the representative of text revision models. As the patent is a form of legal document, legal-specific LLMs may be useful in this task. Hence, we evaluate SaulLM-7B, a model designed for the legal domain (Colombo et al., 2024). For general open-source LLMs, we opt for two types of models with different structures. We include Mixtral-8$\times$7B (Jiang et al., 2024a) based on Sparse Mixture of Experts (SMoE) and the recent Llama-3.1 series (Dubey et al., 2024). We evaluate both the Llama-3.1-8B and Llama-3.1-70B versions to explore the size effect. In addition, we use LoRA (Hu et al., 2021) to fine-tune our own model, Llama-3.1-8B-FT and SaulLM-7B-FT, to investigate the effectiveness of fine-tuning. Fine-tuning details are introduced in Appendix C. We also test the powerful GPT series for comparison, including GPT-3.5 and state-of-the-art GPT-4 (OpenAI, 2023).
@@ 4.2 Evaluation
Human Evaluation To enhance the precision of evaluation for this new task, we incorporate evaluations by patent professionals, adhering strictly to established examination criteria. Recent research (Jiang et al., 2024b) suggests five criteria for assessing the quality of patent claims, which align perfectly with our patent claim revision objectives introduced in the Introduction section.
@@ 5.4 Advantages of Fine-tuning
Table 3 illustrates that fine-tuned Llama-3.1-8B achieves the highest SARI (55.5), while fine-tuned SaulLM-7B reaches the best BLEU (0.63) and R-L (0.67) among all LLMs in automated evaluations. In human evaluation, the fine-tuned models outperform their corresponding base models and also the copy baseline in all aspects, demonstrating the effectiveness of fine-tuning. This finding aligns with previous research on patent claim generation (Jiang et al., 2024b). Particularly, SaulLM-7B-FT achieves almost the same overall quality score (6.38) as GPT-4 (6.40), and it even outperforms GPT-4 in clarity and consistency. This finding suggests that in-domain training would bring significant advantages to patent text generation, a specific type of text featuring high precision.
@@ 5.6 Inconsistency between Automated and Human Evaluations
Takeaways The inconsistency of the automated metrics with the human gold standard shows the limitations of the metrics. G-Eval is a currently more suitable choice to automatically evaluate patent claims. There is still a need for better automated evaluation methods for patents that have closer alignment to human expert evaluation.
@@ B.1 Calculation Method
(a) Number of draft claims and published claims
(b) Number of tokens of draft claims and published claims
Figure 3: Frequency diagram of number of claims and tokens
@@ Appendix D Model Details
Mixtral We also include Mixtral-8$\times$7B that is based on Sparse Mixture of Experts (SMoE) and uses the same architecture as Mistral-7B (Jiang et al., 2024a). We select Mixtral-8$\times$7B-Instruct for less computational costs.
@@
GPT-3.5 We also include GPT series, GPT-3.5, for comparison. Specifically, we use the latest GPT-3.5 Turbo model444gpt-3.5-turbo-0125: https://platform.openai.com/docs/models/gpt-3-5-turbo, which achieves higher accuracy in adhering to specified output formats. This model extends the context window to 16,385 tokens, which supports more examples in the prompt. Similar to the above, we evaluate GPT-3.5 with one-shot prompting, where one example is randomly chosen for each test input.
GPT-4 GPT-4 represents the state-of-the-art LLM with expansive general knowledge and enhanced reasoning capabilities optimized for chat. This capability allows GPT-4 to tackle more complex challenges with increased accuracy (OpenAI, 2023). We use the recent GPT-4 model555gpt-4-0125-preview: https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo, designed to address the problem that the model sometimes does not complete a task. GPT-4 significantly expands the context length to 128,000 tokens, so we use the same experimental setting as GPT-3.5 for fair comparison.
@@ E.2 Standard Automated Metrics for Text Revision
For automated evaluation metrics, we use the package from the HuggingFace evaluate library.

### v2

@@ 1 Introduction
Text revision aims to improve text quality, such as fixing grammar errors (Fang et al., 2023) and enhancing sentence coherence (Geva et al., 2019). Currently, datasets for this task are derived from scientific literature, Wikipedia entries, and news articles (Du et al., 2022). In this paper, we broaden the scope of text revision to encompass the domain of patents, characterized by large-scale, complex, and precise textual data. The patent domain presents unique opportunities and challenges for the field of artificial intelligence (AI) and natural language processing (NLP) (Jiang and Goetz, 2025).
@@
Patent claims are critical in a patent application document. As the legal centerpiece, they define the technical scope of the invention and ensure the patent can withstand legal scrutiny. We introduce the relevant background information of patents in Appendix A. Drafting and revising patent applications are both time-intensive and financially burdensome (LLP, 2023). Research showed that large language models (LLMs) have the potential to generate high-quality patent claims but current performance is not yet satisfactory (Jiang et al., 2025c, a). To facilitate the automation of patent writing, we propose a new task, namely patent claim revision, aiming at improving the quality of patent claims to pass the legal scrutiny of patent offices.
Figure 1 illustrates an example of patent claim revision. By checking the dataset and consulting patent professionals, we have identified five different types of modifications between the draft and final versions. (1) Content amendment: Essential information missing in the draft is included to highlight novel features of the invention, and unnecessary or redundant information is removed, including overlaps with the prior art. (2) Term consistency: Technical terms are ensured to be consistent throughout the document. (3) Language precision: Grammatical errors are corrected, and word choice is refined for greater precision. (4) Concision: Some claims are merged with others for concision. (5) Renumbering: The consolidation of claims necessitates adjustments in their numbering.
@@ 2.2 Patent Writing
The adoption of LLMs in generating patent text primarily aims to enhance the efficiency and efficacy of drafting patent applications. Despite the potential capabilities of LLMs, current research in this area remains limited and largely unsatisfactory (Jiang and Goetz, 2025). An early study by Lee and Hsiang (2020) served as a preliminary exploration into generating patent claims with the fine-tuning of GPT-2 (Radford et al., 2019). The authors demonstrated that minimal training steps were adequate for the model to generate patent-like texts, but they did not evaluate the quality of the generated text. Subsequent research by Lee (2020) expanded on this aspect by training GPT-2 to convert one element of a patent application into another, for example, creating abstracts from titles and claims from abstracts. As the abstract is typically rather generic and imprecise, the latter may not be a well-conditioned task. Hence, Jiang et al. (2025c) proposed the description-based claim generation task and evaluated the performance of the current LLMs on this domain-specific task. We extend the task to claim revision to explore whether LLMs can further improve the quality of claims. Moreover, Christofidellis et al. (2022) introduced a prompt-based generative transformer (PGT) for patent-related tasks, which used GPT-2 as a foundational model and employed multi-task learning (MTL) (Maurer et al., 2016) to train on various tasks, including part-of-patent generation, text infilling, and evaluating patent coherence. Additionally, Aubakirova et al. (2023) presented the first large-scale patent figure-caption dataset, designed for patent figure caption generation.
@@ 3.1 Data Collection
Step 2: The European Patent Office provides the Open Patent Services (OPS) for public access to their data. We retrieve the application and published versions of claims corresponding to specific patent publication numbers through the OPS API. A patent has different versions published by EPO, where A1 or A2 is the patent application and B1 is the granted patent. We primarily used the A1 version, and we used A2 if A1 is not available. We eliminated the patents that do not have either A1 or A2. We opted for the B1 version as the revised claims and discarded those without the B1 version.
Step 3: In the final compilation of our dataset, we formulate data into an easy-readable format and manually check in detail to ensure the dataset’s quality.
@@ 4.1 Models
We use the original claims without any editing as a baseline similar to previous text revision works (Raheja et al., 2023; Jourdan et al., 2024), namely the Copy baseline. This baseline performs well when there are extensive overlaps between source inputs and target outputs. We can also evaluate other models’ performance by comparing the results with this simple baseline. We select the state-of-the-art CoEdIT-XL (Raheja et al., 2023) as the representative of text revision models. As the patent is a form of legal document, legal-specific LLMs may be useful in this task. Hence, we evaluate SaulLM-7B, a model designed for the legal domain (Colombo et al., 2024). For general open-source LLMs, we opt for two types of models with different structures. We include Mixtral-8$\times$7B (Jiang et al., 2024) based on Sparse Mixture of Experts (SMoE) and the recent Llama-3.1 series (Dubey et al., 2024). We evaluate both the Llama-3.1-8B and Llama-3.1-70B versions to explore the size effect. In addition, we use LoRA (Hu et al., 2021) to fine-tune our own model, Llama-3.1-8B-FT and SaulLM-7B-FT, to investigate the effectiveness of fine-tuning. Fine-tuning details are introduced in Appendix C. We also test the powerful GPT series for comparison, including GPT-3.5 and state-of-the-art GPT-4 (OpenAI, 2023).
@@ 4.2 Evaluation
Human Evaluation To enhance the precision of evaluation for this new task, we incorporate evaluations by patent professionals, adhering strictly to established examination criteria. Recent research (Jiang et al., 2025c) suggests five criteria for assessing the quality of patent claims, which align perfectly with our patent claim revision objectives introduced in the Introduction section.
@@ 5.4 Advantages of Fine-tuning
Table 3 illustrates that fine-tuned Llama-3.1-8B achieves the highest SARI (55.5), while fine-tuned SaulLM-7B reaches the best BLEU (0.63) and R-L (0.67) among all LLMs in automated evaluations. In human evaluation, the fine-tuned models outperform their corresponding base models and also the copy baseline in all aspects, demonstrating the effectiveness of fine-tuning. This finding aligns with previous research on patent claim generation (Jiang et al., 2025c). Particularly, SaulLM-7B-FT achieves almost the same overall quality score (6.38) as GPT-4 (6.40), and it even outperforms GPT-4 in clarity and consistency. This finding suggests that in-domain training would bring significant advantages to patent text generation, a specific type of text featuring high precision.
@@ 5.6 Inconsistency between Automated and Human Evaluations
Takeaways The inconsistency of the automated metrics with the human gold standard shows the limitations of the metrics. G-Eval is a currently more suitable choice to automatically evaluate patent claims. There is still a need for better automated evaluation methods for patents that have closer alignment to human expert evaluation. A follow-up work investigated a new evaluation method specifically designed for patent claims (Jiang et al., 2025b).
@@ Appendix A Patent Background
(a) Number of draft claims and published claims
(b) Number of tokens of draft claims and published claims
Figure 3: Frequency diagram of number of claims and tokens
@@ Appendix D Model Details
Mixtral We also include Mixtral-8$\times$7B that is based on Sparse Mixture of Experts (SMoE) and uses the same architecture as Mistral-7B (Jiang et al., 2024). We select Mixtral-8$\times$7B-Instruct for less computational costs.
@@
GPT-3.5 We also include GPT series, GPT-3.5, for comparison. Specifically, we use the latest GPT-3.5 Turbo model333gpt-3.5-turbo-0125: https://platform.openai.com/docs/models/gpt-3-5-turbo, which achieves higher accuracy in adhering to specified output formats. This model extends the context window to 16,385 tokens, which supports more examples in the prompt. Similar to the above, we evaluate GPT-3.5 with one-shot prompting, where one example is randomly chosen for each test input.
GPT-4 GPT-4 represents the state-of-the-art LLM with expansive general knowledge and enhanced reasoning capabilities optimized for chat. This capability allows GPT-4 to tackle more complex challenges with increased accuracy (OpenAI, 2023). We use the recent GPT-4 model444gpt-4-0125-preview: https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo, designed to address the problem that the model sometimes does not complete a task. GPT-4 significantly expands the context length to 128,000 tokens, so we use the same experimental setting as GPT-3.5 for fair comparison.
@@ E.2 Standard Automated Metrics for Text Revision
For automated evaluation metrics, we use the package from the HuggingFace evaluate library. [^5] https://github.com/huggingface/evaluate
@@ Appendix F More Results
Appendix G Broader Impact
The new task of patent claim revision and related challenges provide valuable insights for the broader field of text revision. Current text-revision models cannot perform well on patent claim revision, because of the unique characteristics of patent texts and goals of revision. This task offers a compelling case for further research into different types and domain-specific text revision methodologies. In addition, the results demonstrate the importance of fine-tuning and domain adaptation for this patent task. Compared to specified LLMs, general-purpose LLMs show limited effectiveness, a finding that may apply to other related fields such as legal documentation. Furthermore, the results underscore the limitations of current automated evaluation metrics in structured, domain-specific contexts. The gap between human assessments and automated metrics highlights the need for more robust evaluation methods that better align with human judgment in specialized contexts.